# Stochastic gradient descent estimation of generalized matrix factorization models with application to single-cell RNA sequencing data

### v1

@@ 6.2 TENxBrainData
To demonstrate the scalability of our method to large datasets, we apply SGD to the TENxBrainData (Lun and Morgan,, 2023), which consists of scRNA-seq read counts from approximately 1.3 million cells from the mouse brain, generated by 10X Genomics. Quality control and filtering are performed using the R package scater (McCarthy et al.,, 2017), excluding cells with an exceptional number of mitochondrial reads (more than 3 median absolute deviances away from the median), and genes with no expression in over 99% of the cells. This procedure yields $1{,}232{,}055$ cells and we retain the 500 most variable genes, as done with the Arigoni dataset.
Considering the eigenvalue gap method (Supplementary Fig. 2), which is a very fast procedure for model selection, we selected a model with 10 latent factors, which we were able to fit in 77 minutes with SGD. Subsequent Leiden clustering of this matrix factorization revealed sensible clusters that align with established marker genes (Fig. 5A). For example, cluster 8 is characterized by cells expressing Pyramidal neuron cells marker genes, such as Crym (Loo et al.,, 2019), while cluster 4 contains cells expressing interneurons markers, e.g., Sst and Lhx6 (Tasic et al.,, 2018).
Figure 5: Results on large-scale data. A) Heatmap of the average gene expression of the 18 clusters obtained by Leiden clustering computed on the latent score matrix, for 29 marker genes. Each marker gene is coloured based on the celltype it is expressed in. B) Computational time for different methods on increasingly larger subsets of the dataset. Each subset includes 500 high-variable genes and a growing number of cells.
To study the scalability of SGD and its competitors on this real dataset, we considered subsamples of $100{,}000$, $200{,}000$ and $300{,}000$ cells (Fig. 5B). This analysis showed that SGD is orders of magnitude faster than competing methods, with Fisher and NBWaVE taking $3.5$ and $10$ hours, respectively, to analyze $300{,}000$ cells. While AvaGrad achieves better computational speed than Fisher and NBWaVE, it still takes more for it to analyze $300{,}000$ cells than for SGD to fit the whole 1.2 million cells. Moreover, AvaGrad did not converge one in five times for $100{,}000$ cells, and two in five times for $200{,}000$ cells.
@@ 7 Discussion
In the present work, we propose a flexible and scalable tool to perform generalized matrix factorization in massive data problems, with a particular focus on scRNA-seq applications. We propose an innovative adaptive stochastic gradient descent algorithm, whose performances are enhanced via a memory-efficient block-wise subsampling method and a convenient initialization strategy. An R/C++ implementation in the open-source package sgdGMF is freely available on GitHub. Overall, the proposed method proved competitive with state-of-the-art approaches available in R, showing higher prediction accuracy, good biological signal-extraction ability, and a significant speed-up of the execution time in simulated and real data examples. Unlike most methods currently employed for scRNA-seq signal extraction, our approach natively deals with missing values, iteratively imputing them with the model’s current best prediction. This feature proved important for out-of-sample error evaluation, model selection, and matrix completion.
@@
From an algorithmic viewpoint, another fascinating possibility is considering non-uniform sampling schemes for the mini-batch selection. For instance, if the sample is divided into known subpopulations, it could be convenient to exploit the clustered nature of the data when forming the mini-batch partition via stratification. This can improve the representativeness of each chunk, reduce the variance of the gradient estimator and prevent the optimization from staking on suboptimal maxima dominated by a specific subpopulation signal.
@@ 8 Funding
This work was supported by EU funding within the MUR PNRR “National Center for HPC, big data and quantum computing” (Project no. CN00000013 CN1). DR was also supported by by the National Cancer Institute of the National Institutes of Health (U24CA289073). The views and opinions expressed are only those of the authors and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. This work was supported by grants from Ghent University Special Research Fund (BOF20/GOA/023) (A.S., L.C.), Research Foundation Flanders (FWO G062219N) (A.S., L.C.).
@@ 9 Data availability
sgdGMF is freely available as an open-source R package at github/CristianCastiglione/sgdGMF. The scripts used to run all analyses are available on GitHub at github/alexandresegers/sgdGMF_Paper. The Arigoni dataset is available at https://doi.org/10.6084/m9.figshare.23939481.v1. The TENxBrainData dataset is available as part of the TENxBrainData Bioconductor package at https://bioconductor.org/packages/TENxBrainData.

### v2

@@ 6.2 TENxBrainData
To demonstrate the scalability of our method to large datasets, we apply aSGD to the TENxBrainData (Lun and Morgan,, 2023), which consists of scRNA-seq UMI counts generated by 10X Genomics for approximately 1.3 million cells. Cells of this dataset were obtained from the cortex, hippocampus and ventricular zone of two mouse brains. Although no ground truth is available for the different cell types in this dataset, marker genes that can discriminate between subtypes of cells are available for mouse brains, and the list used by Hicks et al., (2021) is used to qualitatively evaluate extraction of biological signal in different cell clusters. Quality control and filtering are performed using the R package scater (McCarthy et al.,, 2017), excluding cells with an exceptional number of mitochondrial reads (more than 3 median absolute deviances away from the median), and genes with no expression in over 99% of the cells. This procedure yields 1,232,0551{,}232{,}055 cells and we retain the 500 most variable genes, as done with the Arigoni dataset.
Considering the eigenvalue gap method (Supplementary Fig. 4), which is a very fast procedure for model selection, we selected a model with 10 latent factors. To study the scalability of aSGD and its competitors on this large dataset, we considered subsamples of 100,000100{,}000, 200,000200{,}000 and 300,000300{,}000 cells (Fig. 6A). This analysis showed that aSGD is orders of magnitude faster than competing methods, with Fisher and NBWaVE taking $4$ and $8$ hours, respectively, to analyze 300,000300{,}000 cells. While AvaGrad achieves better computational speed than Fisher and NBWaVE, it remains extremely slow compared to aSGD. Moreover, AvaGrad did not converge one in five times for 100,000100{,}000 cells, and two in five times for 200,000200{,}000 cells. Therefore, these methods would result in extreme computational times on the full dataset. Moreover, the memory usage of aSGD remains lower than competing methods. Note also that COAP returned errors when using more than 70,00070{,}000 cells, rendering it unable to handle large datasets. Therefore, only aSGD is a reasonable model to run on the full dataset due to its superior computational efficiency.
Using aSGD on the full dataset returned results in 77 minutes. Subsequent Leiden clustering of this matrix factorization revealed sensible biological signal extraction, as its clusters align with established marker genes (Fig. 6B). For example, cluster 8 is characterized by cells expressing Pyramidal neuron cells marker genes, such as Crym (Loo et al.,, 2019), while cluster 4 contains cells expressing interneuron markers, e.g., Sst and Lhx6 (Tasic et al.,, 2018).
Figure 6: Results on large-scale data. A) Computational time and memory usage for different methods on increasingly larger subsets of the dataset. Each subset includes 500 high-variable genes and a growing number of cells. B) Heatmap of the average gene expression of the 18 clusters obtained by Leiden clustering computed on the latent score matrix, for 29 marker genes. Each marker gene is coloured based on the cell type it is expressed in.
@@ 7 Discussion
In the present work, we propose a flexible and scalable tool to perform generalized matrix factorization in massive data problems, with a particular focus on scRNA-seq applications. We propose an innovative adaptive stochastic gradient descent algorithm, whose performances are enhanced via a memory-efficient block-wise subsampling method and a convenient initialization strategy. An R/C++ implementation in the open-source package sgdGMF is freely available on CRAN. Overall, the proposed method proved competitive with state-of-the-art approaches available in R, showing higher prediction accuracy, good biological signal-extraction ability, and a significant speed-up of the execution time in simulated and real data examples. Unlike most methods currently employed for scRNA-seq signal extraction, our approach natively deals with missing values, iteratively imputing them with the model’s current best prediction. This feature proved important for out-of-sample error evaluation, model selection, and matrix completion.
@@
From an algorithmic viewpoint, another fascinating possibility is considering non-uniform sampling schemes for the mini-batch selection. For instance, if the sample is divided into known subpopulations, it could be convenient to exploit the clustered nature of the data when forming the mini-batch partition via stratification. This can improve the representativeness of each chunk, reduce the variance of the gradient estimator, and prevent the optimization from converging to suboptimal maxima dominated by a specific subpopulation signal.
@@ 8 Funding
This work was supported by EU funding within the MUR PNRR “National Center for HPC, big data and quantum computing” (Project no. CN00000013 CN1). The views and opinions expressed are only those of the authors and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them. DR was also supported by the National Cancer Institute of the National Institutes of Health (U24CA289073) and by project EOSS6-0000000644 from the Chan Zuckerberg Initiative. This work was supported by grants from Ghent University Special Research Fund (BOF20/GOA/023) (A.S., L.C.), Research Foundation Flanders (FWO G062219N) (A.S., L.C.).
@@ 9 Data availability
sgdGMF is freely available as an open-source R package at CRAN.R-project.org/package=sgdGMF. The scripts used to run all analyses are available on GitHub at github/alexandresegers/sgdGMF_Paper. The Arigoni dataset is available at https://doi.org/10.6084/m9.figshare.23939481.v1. The TENxBrainData dataset is available as part of the TENxBrainData Bioconductor package at https://bioconductor.org/packages/TENxBrainData.